OpenStudy (anonymous):

Write the equation of the line that is parallel to the line y = 2x + 2 and passes through the point (5, 3). y = 2x − 7 y = −x − 2 y = 2x − 2 y = −x − 7

OpenStudy (anonymous):

State the slope which in this case is 2 according to y=mx+b

OpenStudy (anonymous):

Then plug in y=mx+b and 2 as m and then plug in (5, 3) as x and y. Then solve for b.

OpenStudy (anonymous):

so its either choice A or C

OpenStudy (anonymous):

Correct. Because the slope has to be 2 because parallel lines have the same slope.

OpenStudy (anonymous):

im pretty sure its A am i correct?

OpenStudy (anonymous):

Well, lets see. y=mx+b we got 2 as m and we have the coordinate (5,3) x=5 y=3 Plug it in the formula. y=mx+b 3=2(5)+b 3=10+b subtract 10 from both sides. -7=b So now we got all neccesary items to rewrite this as another equation thats parallel to the previous equation. We got the slope: 2 We got the b: -7 Plug it all in. y=mx+b y=2x-7

OpenStudy (anonymous):

oh ok thanks alot!

OpenStudy (anonymous):

Sure no problem :D